From GlueXWiki
Jump to: navigation, search

May 5, 2011 FDC meeting


  1. Production
    • Status (Dave)
    • Wire position scanning logbook page 564 (Lubomir)
    • End window material (Bill)
  2. Electronics (Fernando, Chris)
    • Status
    • Rigid-flex stuffing
  3. Engineering status (Bill)
  4. Chamber testing (Beni)
  5. Other


Participants: Eugene, Bill, Dave, Roger, Beni, and Lubomir.


- Dave: Two cathode frames type 3 (with ground planes with holes) are assembled waiting for the rigid-flexes and ground wires to be glued on them. One wire frame was strung, the wires glued and now being soldered. Expect to be ready by the end of the day, then we will do position and tension measurements. We started also cutting foils for the third cathode plane and the second wire frame is also ready.

- We discussed the configuration needed to test the first chamber: end window, cathode type 3, wire plane, cathode type 3, end window, plus empty frames to make the same thickness as of the full package. We will use the plastic gussets. Bill wanted first to do gas leakage tests with these plastic gussets, but actually such test was done before.

- End window material. Bill found two options: 1mil available now, and 2mil roll for ~$1500 total that we can have in ~2 weeks. For the first two end windows we decided to use the 1mil material.

- Wire position measurements. We had problems finding the field wire positions: the laser sensor doesn't see about 20% of them while sees all the sense wires. Possible explanations: vibrations due to HVAC on the top of the clean room; one can actually see the weights vibrating. It could be also the surface of the field wires is somehow different in reflecting the laser beam. Before for the test wire frame all the field wires were seen by the sensor, but we used different type of field wires. Two samples of our field wires will be analyzed with electron microscope by Olga. It turned out that if the sensor doesn't see the wire, a tiny move along the wire is enough to see it. This is the way all the wires of the first chamber were scanned before gluing them. These results are compared with the old results for the test chamber at page 564 of the logbook linked above. In between the two stringing the jig rail was modified to fix the ~250microns offset in one pitch, but now it looks like we have ~80 microns in the opposite direction. Eugene also mentioned the difference in the slope in the above plots before and after for which we couldn't find an explanation.

- We were able to fix all the sense wires that were off by more than ~60 microns (6 wires in total) by re-taping them. Also right after gluing we scanned all the sense wires. It took ~1 hour and in principle you could correct the wire positions if you have to, since the glue is still not hard enough.


- Lubomir talked to Fernando about assembling by Chris the rigid-flexes that are needed for the first chamber. If we can have the assembled flexes back from the company within a week that would be fine, but if not we should consider doing it here. We should decide about this by next Monday. Also there was a question to Fernando: what kind of tests have to be done with the assembled rigid-flexes before gluing them to the cathodes.

- Roger brought the second batch of the rigid-flexes, to be moved to 117.


- Bill: the ultrasonic cleaner that was fixed by the manufacturer will be back next week. Bill was looking also for a storage for the produced chambers.

Chamber testing

- The parts needed for the chamber testing at Blue Crab are listed above by Beni. Mark will help him with the gas connectors. We have now only one Bertan supply that we have to share between 126 and Blue Crab. The second supply has to be fixed but they asked for $1700 just to look at it; Mark will try to fix it. We need also blind penetration in the clean room walls to have gas inside. Eugene will try to expedite the work order made weeks ago.

- Beni connected a pressure sensor to measure the pressure difference between the top and middle chamber of the prototype package, but the readings fluctuate because of the bubbling. The best is to connect the sensor to the extra tubes on the spacer rings, but those are broken and the fix them most likely will require disassembly of that package.